OpenStudy (anonymous):

Help Please!! Solve this equation by completing the square. 3x^2 - 4x = -2

OpenStudy (gorv):

make coefficient of x^2 unity x^2-4/3x=-2/3

OpenStudy (gorv):

x^2-2*x*2/3+2/3=0

OpenStudy (gorv):

x^2-2*x*2/3+(2/3)^2-(2/3)^2+2/3=0

OpenStudy (gorv):

x^2-2*x*2/3+(2/3)^2=(x-2/3)^2 (x-2/3)^2-4/9+2/3=0 (x-2/3)^2-2/9=0 (x-2/3)^2=2/9 x-2/3=(2/9)^1/2 solve now

OpenStudy (anonymous):

where did the ^1/2 come from tho?

OpenStudy (gorv):

(x-2/3)^2=2/9 take square root of both sides